OpenStudy (anonymous):

Solve for x. -5x − 4 = -29 A. x = 5.6 B. x = -5 C. x = 5 D. x = 4

OpenStudy (bahrom7893):

add 4 to both sides. what do u get?

OpenStudy (anonymous):

both side of what?

OpenStudy (bahrom7893):

of the equation

OpenStudy (anonymous):

-9 and 8? .-.

OpenStudy (bahrom7893):

Well: -5x-4 + 4 = -29 + 4

OpenStudy (bahrom7893):

What does that simplify to?

OpenStudy (uri):

-5x − 4 = -29 +4 +4 ========= @renae_ Solve this.

OpenStudy (anonymous):

-5 = -25 ????

OpenStudy (bahrom7893):

Close: -5x = -25

OpenStudy (bahrom7893):

Now divide both sides by -5. What do you get?

OpenStudy (anonymous):

-1x = -5 ??

OpenStudy (uri):

-5x=-25*

OpenStudy (bahrom7893):

well that's if you divide by 5. So if -1x = -5, what does x equal to?

OpenStudy (bahrom7893):

-x=-5 x = ?

OpenStudy (uri):

You can do it this way Divide -25/-5.. x=-25/-5 Divide and answer.

OpenStudy (anonymous):

C.? 5?

OpenStudy (bahrom7893):

yes

OpenStudy (anonymous):

yay thanks (:

OpenStudy (uri):

Yes,always remember double negative=positive.